# Methods and Trends in Detecting Generated Images: A Comprehensive Review

### v1

@@ 2.2.8. Representation from Encoder-Decoder for Image Detection by Koutlis: 2025
Table 2. Evaluation of Detection Models from Pretrained Vision-Language Methods Category on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization
@@
Ojha et al. (Ojha et al., 2023) X
Wu et al. (Wu et al., 2023) X
Zhu et al. (Zhu et al., 2023) X
Liu et al. (Liu et al., 2024b) X
Cao et al. (Cao et al., 2024) X
Liu et al. (Liu et al., 2024a) X
Cozzolino et al. (Cozzolino et al., 2024) X
Koutlis et al. (Koutlis and Papadopoulos, 2025) X
Table 3. Evaluation of multimodal vision-language methods on GAN and diffusion-generated images using the UnivFD dataset (Ojha et al., 2023). Results are reported in classification accuracy (%). Methods from (Cozzolino et al., 2024) and (Koutlis and Papadopoulos, 2025) were trained on only four classes of ProGAN-generated images, while others were trained on the full ProGAN dataset from (Wang et al., 2020b).
@@
Ojha et al. (Ojha et al., 2023) 100.0 98.50 94.50 82.00 99.50 97.00 66.60 63.00 57.50 59.5 72.00 70.03 94.19 73.76 94.36 79.07 79.85 78.14 86.78 81.38
Zhu et al. (Zhu et al., 2023) 99.00 99.50 99.30 99.05 99.00 96.75 88.20 63.50 67.50 93.90 98.75 98.70 98.80 98.60 98.75 98.75 98.75 98.75 98.45 94.42
Liu et al. (Liu et al., 2024b) 100.0 99.33 99.67 99.46 99.83 97.07 77.53 81.11 65.50 82.32 96.79 90.70 98.30 95.90 98.75 92.40 93.95 93.00 94.90 92.45
Cao et al. (Cao et al., 2024) 100.0 97.40 97.50 97.50 96.20 98.65 73.85 93.00 75.00 92.75 93.20 77.35 98.70 96.60 98.80 87.75 89.95 88.70 97.00 92.10
Liu et al. (Liu et al., 2024a) 99.90 99.30 99.50 97.20 99.40 99.80 93.20 _ _ _ _ 76.10 98.60 94.90 98.70 94.40 94.70 94.20 98.80 _
Koutlis et al. (Koutlis and Papadopoulos, 2025) 100.0 99.30 99.60 88.90 99.80 99.50 80.60 90.60 68.30 89.20 90.60 76.10 98.30 88.20 98.60 88.90 92.60 90.70 95.00 _
2.3. Frequency Domain Analysis Methods
@@
Frequency domain analysis transforms image data into the spectral domain, facilitating the detection of periodic artifacts, noise distributions, and variations in frequency components often associated with synthetic image generation. Techniques such as the Discrete Fourier Transform (DFT), Discrete Wavelet Transform (DWT), and Discrete Cosine Transform (DCT) are commonly used to extract these spectral features. The Fourier Transform, proposed by Fourier (Baron Fourier et al., 2003), decomposes signals into their constituent frequencies and is effective in identifying global periodic patterns, including checkerboard artifacts. The two-dimensional Discrete Fourier Transform (DFT) of an image $f(x,y)$ of size $M\times N$ is given by:
(5) $F(u,v)=\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}f(x,y)e^{-j2\pi\left(\frac{ux}{M}+\frac{vy}{N}\right)}$
@@
(6) $f(x,y)=\frac{1}{MN}\sum_{u=0}^{M-1}\sum_{v=0}^{N-1}F(u,v)e^{j2\pi\left(\frac{ux}{M}+\frac{vy}{N}\right)}$
To improve computational efficiency, the Fast Fourier Transform (FFT), introduced by Cooley and Tukey (Cooley and Tukey, 1965), accelerates the DFT process, enabling the rapid detection of periodic patterns and aliasing artifacts commonly found in synthetic images. The Wavelet Transform, introduced by Haar (Haar, 1911), provides both spatial and frequency localization, making it suitable for detecting transient and localized artifacts. Unlike the Fourier Transform, which represents signals in terms of sinusoidal waves, the Discrete Wavelet Transform (DWT) uses wavelet basis functions to analyze variations in different frequency bands. The one-level decomposition of an image using DWT is given by:
(7) $f(x,y)=\sum_{m}\sum_{n}c_{m,n}\phi_{m,n}(x,y)+\sum_{m}\sum_{n}d_{m,n}\psi_{m,n}(x,y)$
where $\phi_{m,n}(x,y)$ represents the approximation coefficients (low-frequency components), and $\psi_{m,n}(x,y)$ captures the detailed coefficients (high-frequency components) at different scales. The Discrete Cosine Transform (DCT), widely applied in image compression (e.g., JPEG), is particularly effective for analyzing energy distribution in smooth regions and identifying compression-induced anomalies (Ahmed et al., 1974). The 2D DCT of an image block $f(x,y)$ of size $N\times N$ is computed as:
(8) $F(u,v)=\alpha(u)\alpha(v)\sum_{x=0}^{N-1}\sum_{y=0}^{N-1}f(x,y)\cos\left[\frac{(2x+1)u\pi}{2N}\right]\cos\left[\frac{(2y+1)v\pi}{2N}\right]$
@@
(9) $\alpha(u)=\begin{cases}\frac{1}{\sqrt{N}},&\text{if }u=0\\ \frac{\sqrt{2}}{\sqrt{N}},&\text{if }u>0\end{cases}$
@@
Figure 4. Comparison of frequency domain transformations for real and generated images. The first column presents the input images, followed by their respective DFT, DWT, and DCT representations. The images were obtained from ForenSynth (Wang et al., 2020b), and the generated image was produced using CycleGAN (Zhu et al., 2017). These images are reproduced under the Attribution-NonCommercial-ShareAlike 4.0 International license.
[…]
Diagram illustrating different frequency transformation methods.
2.3.1. Simulation Artifacts and Detection in Frequency Domain by Zhang: 2019
Zhang et al. (Zhang et al., 2019) proposed a detection framework leveraging frequency-domain analysis to identify GAN-generated images. Their key contributions include the introduction of AutoGAN, a GAN simulator that mimics artifacts found in diverse GAN architectures, and the use of frequency spectrum inputs instead of spatial features for classification. AutoGAN generates synthetic images by simulating the up-sampling artifacts present in GAN-generated images, independent of specific GAN architectures. The generator applies adversarial loss as shown in Table 1 and an $l_{1}$-norm loss function to minimize differences between generated and real images. This allows training without requiring access to actual GAN-generated images. To extract meaningful features in the frequency domain, the framework applies a 2D Discrete Fourier Transform (DFT), as given by equation (5), to RGB images, generating three-channel frequency representations. These representations undergo logarithmic transformation and normalization within the range before being used for classification. For classification, a ResNet-34 model, pre-trained on ImageNet (Russakovsky et al., 2015), processes the transformed frequency features.
2.3.2. Fourier Spectrum Discrepancies: Dzanic 2020
@@
Dzanic et al. (Dzanic et al., 2020) presented a method for analyzing high-frequency Fourier models to highlight the limitations of generative models, such as GANs and VAEs, in reconstructing high-frequency components. Their approach involves applying the Fourier Transform to images to obtain a reduced spectrum, which is then modeled using two decay parameters: $b_{1}$, representing the magnitude of high-frequency content, and $b_{2}$, representing the decay rate. These parameters were used to train a KNN classifier capable of distinguishing synthetic images from real ones, achieving 99.2% accuracy on uncompressed high-resolution images. The process involves normalizing the Fourier transform by the DC gain, converting the data to normalized polar coordinates, binning and averaging Fourier coefficients to create a reduced spectrum, and fitting the decay parameters above a threshold wavenumber. This comprehensive method underscores the effectiveness of frequency domain features in identifying the discrepancies characteristic of synthetic image generation.
2.3.3. Liu’s Detection Method Derived from Analysis on Real Images: 2022
@@
Liu et al. (Liu et al., 2022b) proposed a novel approach to detecting synthetic images by focusing on the inherent noise patterns of real images, deviating from existing methods that analyze artifacts in generated images. They introduced the concept of Learned Noise Patterns (LNP), a high-dimensional spatial mapping derived from neural networks, to characterize the noise properties of real images. By comparing these learned patterns with the noise present in synthetic images, the method identifies discrepancies that indicate image generation. Leveraging both spatial and frequency domain representations, this approach demonstrated improved accuracy in detecting synthetic images across multiple domains.
2.3.4. Two-Stream Convolutional Network for Fake Content Detection by Yousaf: 2022

### v2

@@ 2.2.11 ThinkFake: Reasoning in Multimodal Large Language Models by Huang et al: 2025
Table 2: Evaluation of Detection Models from Pretrained Vision-Language Methods (and/or Multi-modal Large Language Model) Category on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization
@@
Ojha et al. Ojha et al. (2023) X
Wu et al. Wu et al. (2023) X
Zhu et al. Zhu et al. (2023) X
Liu et al. Liu et al. (2024b) X
Cao et al. Cao et al. (2024) X
Liu et al. Liu et al. (2024a) X
Cozzolino et al. Cozzolino et al. (2024) X
Koutlis et al. Koutlis and Papadopoulos (2025) X
Ji et al. Ji et al. (2025a) X
Ji et al. Ji et al. (2025b) X
Huang et al. Huang et al. (2025) X
Tan et al. Tan et al. (2025) X
Zhou et al. Zhou et al. (2025)
Fan et al. Fan et al. (2024) X
Table 3: Evaluation of multimodal vision-language methods on GAN and diffusion-generated images using the UnivFD dataset Ojha et al. (2023). Results are reported in classification accuracy (%). Methods from Cozzolino et al. (2024) and Koutlis and Papadopoulos (2025) were trained on only four classes of ProGAN-generated images, while others were trained on the full ProGAN dataset from Wang et al. (2020b).
@@
Ojha et al. Ojha et al. (2023) 100.0 98.50 94.50 82.00 99.50 97.00 66.60 63.00 57.50 59.5 72.00 70.03 94.19 73.76 94.36 79.07 79.85 78.14 86.78 81.38
Zhu et al. Zhu et al. (2023) 99.00 99.50 99.30 99.05 99.00 96.75 88.20 63.50 67.50 93.90 98.75 98.70 98.80 98.60 98.75 98.75 98.75 98.75 98.45 94.42
Liu et al. Liu et al. (2024b) 100.0 99.33 99.67 99.46 99.83 97.07 77.53 81.11 65.50 82.32 96.79 90.70 98.30 95.90 98.75 92.40 93.95 93.00 94.90 92.45
Cao et al. Cao et al. (2024) 100.0 97.40 97.50 97.50 96.20 98.65 73.85 93.00 75.00 92.75 93.20 77.35 98.70 96.60 98.80 87.75 89.95 88.70 97.00 92.10
Liu et al. Liu et al. (2024a) 99.90 99.30 99.50 97.20 99.40 99.80 93.20 _ _ _ _ 76.10 98.60 94.90 98.70 94.40 94.70 94.20 98.80 _
Koutlis et al. Koutlis and Papadopoulos (2025) 100.0 99.30 99.60 88.90 99.80 99.50 80.60 90.60 68.30 89.20 90.60 76.10 98.30 88.20 98.60 88.90 92.60 90.70 95.00 _
2.3 Training-Free Methods
Training-Free Methods refer to detection approaches that do not require any supervised training on synthetic or real datasets. Instead of learning discriminative boundaries through model optimization, these methods rely on intrinsic analytical cues, statistical inconsistencies, or reconstruction behaviors inherent to the image or its representation. They typically exploit properties observable through pretrained encoders, perturbation responses, or entropy measures to distinguish genuine from generated content. By eliminating dependence on dataset-specific training, training-free methods emphasize generalization, interpretability, and robustness to unseen generative models.
2.3.1 AEROBLADE: Training-Free Detection of Latent Diffusion Images Using Autoencoder Reconstruction Error by Ricker et al., 2024
AEROBLADE Ricker et al. (2024) presents a training-free detection framework for identifying images generated by latent diffusion models (LDMs) through their inherent autoencoder (AE) structure. The core idea is that an LDM’s AE reconstructs generated images with lower error than real ones because synthetic samples lie within the learned latent manifold, while real images fall slightly outside it. Consequently, the reconstruction error between an input image and its AE-reconstructed version serves as a reliable indicator of synthesis. This reconstruction-based approach is comparable to DIRE Wang et al. (2023a) but differs fundamentally in that AEROBLADE requires no training or classifier tuning. Formally, AEROBLADE computes the LPIPS Zhang et al. (2018) distance between an image $x$ and its reconstruction $D(E(x))$. To generalize across generators, reconstruction errors are evaluated over multiple AEs, and the minimum reconstruction error is used as the detection criterion. The framework was evaluated on seven LDMs and achieved near-perfect detection performance (mean AP up to 0.999), generalizing effectively even to closed-source generators such as Midjourney MidJourney (2024). Furthermore, it enables qualitative forensics by visualizing reconstruction error maps that reveal inpainted or manipulated regions within real images. The approach remains robust to perturbations such as JPEG compression, cropping, blur, and noise.
2.3.2 RIGID: A Training-Free and Model-Agnostic Framework for Robust AI-Generated Image Detection by He et al., 2024
RIGID He et al. (2024) introduces a training-free and model-agnostic method for detecting AI-generated images by exploiting the sensitivity of image representations to small noise perturbations. The central observation is that real images exhibit greater stability under minor perturbations than AI-generated images in the feature space of large vision models. RIGID compares the cosine similarity between an image’s feature embedding and that of its noise-perturbed version using pretrained extractors such as DINOv2 Oquab et al. (2023). Images showing lower similarity (determined with the threshold) under perturbation are classified as AI-generated. Unlike training-based detectors DIRE Wang et al. (2023a) or training-free detectors such as AEROBLADE Ricker et al. (2024), RIGID operates during inference, requiring no fine-tuning or prior knowledge of the generative model. Its detection score is derived as:
$S(x)=1\{sim(f(x),f(x+\lambda\delta))\leq\epsilon\},\quad\delta\sim N(0,I),$
where $f$ is the feature extractor and $\lambda$ controls perturbation intensity. The framework generalizes across backbones, with DINOv2 achieving the best trade-off between global and local representation. Evaluated on diverse datasets including ImageNet, LSUN-Bedroom, and GenImage—covering diffusion, GAN, and VAE models—RIGID achieved mean AUC/AP improvements exceeding 25% over AEROBLADE and frequently surpassed trained detectors such as DIRE Wang et al. (2023a) and Corvi et al. Corvi et al. (2023).
2.3.3 HFI: A Unified Framework for Training-Free Detection and Implicit Watermarking of Latent Diffusion Model Generated Images by Choi et al., 2024
HFI Choi et al. (2024) proposes a training-free approach that improves the robustness of autoencoder-based detection of latent diffusion model (LDM) images. It addresses the key limitation of AEROBLADE Ricker et al. (2024), which relies on reconstruction error but tends to overfit to background information, failing on images with simple or uniform backgrounds. HFI instead measures the aliasing effect—distortions of high-frequency components introduced by the LDM autoencoder by quantifying how much high-frequency information is lost during reconstruction. HFI computes the directional derivative of the reconstruction distance in the direction of high-frequency components filtered by a low-pass kernel, capturing subtle distortions beyond global reconstruction loss. A numerical approximation via first-order Taylor expansion allows efficient evaluation without training or model fine-tuning. When multiple autoencoders are available, HFI adopts an ensemble strategy that selects the minimum response across them, improving generalization to unknown generators. Compared to prior training-free methods, HFI achieves clear architectural advancement: unlike RIGID He et al. (2024), which relies on feature-space perturbation stability, and AEROBLADE, which measures pixel-space reconstruction consistency, HFI directly exploits the frequency-space degradation intrinsic to LDM autoencoders. This design allows both detection and model attribution (implicit watermarking) by associating distinct aliasing signatures with specific diffusion models. Empirical evaluation across GenImage, SynthBuster, and DiffusionFace benchmarks shows that HFI consistently surpasses AEROBLADE and RIGID, achieving competitive performance to training-based methods.
2.3.4 Zero-Shot Detection of AI-Generated Images by Cozzolino et al., 2024
Zero-Shot Entropy-based Detector (ZED) Cozzolino et al. (2024) introduces a training-free and generator-independent method for detecting AI-generated images, similar in spirit to AEROBLADE Ricker et al. (2024) and RIGID He et al. (2024), but differing in that it requires training only on real images. Instead of relying on reconstruction or perturbation cues, ZED measures the surprise of an image under a probabilistic model of real imagery learned through a lossless image encoder. The key idea is that real images conform to the learned distribution, while synthetic ones deviate from it, resulting in a higher coding cost. Using the Super-Resolution based lossless Compressor (SReC) Cao et al. (2020), ZED predicts conditional probability distributions of pixels across multiple resolutions and computes the negative log-likelihood (NLL) and entropy maps. The difference between the actual and expected coding costs—termed the coding cost gap—serves as the detection statistic. Synthetic images exhibit larger gaps, especially at higher resolutions, reflecting inconsistencies with the statistical model of natural images. This approach generalizes across both GAN- and diffusion-generated images without dependence on any generator-specific artifacts.
Comparative Analysis of Training-Free Methods
Training-free detection methods share the common objective of identifying AI-generated content without requiring supervised learning, instead leveraging analytical properties inherent to pretrained models or statistical characteristics of natural images. Despite this shared foundation, their core mechanisms diverge in various perspectives. AEROBLADE Ricker et al. (2024) pioneers reconstruction-based detection by exploiting the autoencoder behavior in latent diffusion models, demonstrating that generated images exhibit lower reconstruction error than real ones. RIGID He et al. (2024) extends this paradigm by evaluating the stability of image embeddings under random perturbations, introducing feature-space sensitivity as a discriminative signal independent of any generator. HFI Choi et al. (2024) advances the idea further through frequency-domain analysis, identifying aliasing artifacts in high-frequency components of reconstructed images, thus unifying detection and implicit watermarking. ZED Cozzolino et al. (2024) reinterprets detection as a coding-cost analysis problem, employing a probabilistic model of real images to measure deviations in entropy and compression cost, offering a fully zero-shot and generator-agnostic solution. Collectively, these approaches trace a clear methodological evolution—from pixel and latent reconstruction analysis (AEROBLADE), to perturbation and representation stability (RIGID), to spectral and aliasing cues (HFI), and finally to probabilistic learning (ZED). Importantly, these methods converge toward efficient forensic detection without relying on extensive training.
Table 4: Evaluation of Detection Models from Training-Free Methods on Cross-Family Generators, Cross-Category, and Cross-Scene Generalization.
Models Cross-Family Generators Cross-Category Cross-Scene
AEROBLADE Ricker et al. (2024) X X
RIGID He et al. (2024)
HFI Choi et al. (2024) X X
ZED Cozzolino et al. (2024) X X
Table 5: Diffusion-Based Evaluation of Training-Free Methods and an MLLM Method (ForenX Tan et al. (2025)) on the GenImage dataset Zhu et al. (2024). Accuracy (Acc) is reported.
Method ADM BigGAN GLIDE Midjourney SD1.4 SD1.5 VQDM wukong Mean
RIGID He et al. (2024) 0.790 0.976 0.964 0.797 0.698 0.699 0.860 0.708 0.812
AEROBLADE Ricker et al. (2024) 0.838 0.986 0.990 0.988 0.983 0.984 0.723 0.984 0.935
HFI Choi et al. (2024) 0.923 0.996 0.995 0.998 0.998 0.998 0.905 0.999 0.977
ForenX Tan et al. (2025) 0.974 0.978 0.980 0.979 0.978 0.977 0.977 0.980 0.978
Notes: All methods are training-free diffusion image detectors evaluated on the GenImage dataset Zhu et al. (2024), with an additional method from the MLLM category, ForenX Tan et al. (2025). Unlike training-free approaches, ForenX was trained on SDv1.4-generated images and provides reasoning to explain why an image is classified as generated. The results on training-free methods are reported from the study presented by Choi et al. Choi et al. (2024).
2.4 Frequency Domain Analysis Methods
@@
Frequency domain analysis transforms image data into the spectral domain, facilitating the detection of periodic artifacts, noise distributions, and variations in frequency components often associated with synthetic image generation. Techniques such as the Discrete Fourier Transform (DFT), Discrete Wavelet Transform (DWT), and Discrete Cosine Transform (DCT) are commonly used to extract these spectral features. The Fourier Transform, proposed by Fourier Baron Fourier et al. (2003), decomposes signals into their constituent frequencies and is effective in identifying global periodic patterns, including checkerboard artifacts. The two-dimensional Discrete Fourier Transform (DFT) of an image $f(x,y)$ of size $M\times N$ is given by:
$F(u,v)=\sum_{x=0}^{M-1}\sum_{y=0}^{N-1}f(x,y)e^{-j2\pi\left(\frac{ux}{M}+\frac{vy}{N}\right)}$ (1)
@@
$f(x,y)=\frac{1}{MN}\sum_{u=0}^{M-1}\sum_{v=0}^{N-1}F(u,v)e^{j2\pi\left(\frac{ux}{M}+\frac{vy}{N}\right)}$ (2)
To improve computational efficiency, the Fast Fourier Transform (FFT), introduced by Cooley and Tukey Cooley and Tukey (1965), accelerates the DFT process, enabling the rapid detection of periodic patterns and aliasing artifacts commonly found in synthetic images. The Wavelet Transform, introduced by Haar Haar (1911), provides both spatial and frequency localization, making it suitable for detecting transient and localized artifacts. Unlike the Fourier Transform, which represents signals in terms of sinusoidal waves, the Discrete Wavelet Transform (DWT) uses wavelet basis functions to analyze variations in different frequency bands. The one-level decomposition of an image using DWT is given by:
$f(x,y)=\sum_{m}\sum_{n}c_{m,n}\phi_{m,n}(x,y)+\sum_{m}\sum_{n}d_{m,n}\psi_{m,n}(x,y)$ (3)
where $\phi_{m,n}(x,y)$ represents the approximation coefficients (low-frequency components), and $\psi_{m,n}(x,y)$ captures the detailed coefficients (high-frequency components) at different scales. The Discrete Cosine Transform (DCT), widely applied in image compression (e.g., JPEG), is particularly effective for analyzing energy distribution in smooth regions and identifying compression-induced anomalies Ahmed et al. (1974). The 2D DCT of an image block $f(x,y)$ of size $N\times N$ is computed as:
$F(u,v)=\alpha(u)\alpha(v)\sum_{x=0}^{N-1}\sum_{y=0}^{N-1}f(x,y)\cos\left[\frac{(2x+1)u\pi}{2N}\right]\cos\left[\frac{(2y+1)v\pi}{2N}\right]$ (4)
@@
$\alpha(u)=\begin{cases}\frac{1}{\sqrt{N}},&\text{if }u=0\\ \frac{\sqrt{2}}{\sqrt{N}},&\text{if }u>0\end{cases}$ (5)
@@
Figure 4: Comparison of frequency domain transformations for real and generated images. The first column presents the input images, followed by their respective DFT, DWT, and DCT representations. The images were obtained from ForenSynth Wang et al. (2020b), and the generated image was produced using CycleGAN Zhu et al. (2017). These images are reproduced under the Attribution-NonCommercial-ShareAlike 4.0 International license.
2.4.1 Fourier Spectrum Discrepancies: Dzanic 2020
@@
Dzanic et al. Dzanic et al. (2020) presented a method for analyzing high-frequency Fourier models to highlight the limitations of generative models, such as GANs and VAEs, in reconstructing high-frequency components. Their approach involves applying the Fourier Transform to images to obtain a reduced spectrum, which is then modeled using two decay parameters: $b_{1}$, representing the magnitude of high-frequency content, and $b_{2}$, representing the decay rate. These parameters were used to train a KNN classifier capable of distinguishing synthetic images from real ones, achieving 99.2% accuracy on uncompressed high-resolution images. The process involves normalizing the Fourier transform by the DC gain, converting the data to normalized polar coordinates, binning and averaging Fourier coefficients to create a reduced spectrum, and fitting the decay parameters above a threshold wavenumber. This comprehensive method underscores the effectiveness of frequency domain features in identifying the discrepancies characteristic of synthetic image generation.
2.4.2 Liu’s Detection Method Derived from Analysis on Real Images: 2022
@@
Liu et al. Liu et al. (2022b) proposed a novel approach to detecting synthetic images by focusing on the inherent noise patterns of real images, deviating from existing methods that analyze artifacts in generated images. They introduced the concept of Learned Noise Patterns (LNP), a high-dimensional spatial mapping derived from neural networks, to characterize the noise properties of real images. By comparing these learned patterns with the noise present in synthetic images, the method identifies discrepancies that indicate image generation. Leveraging both spatial and frequency domain representations, this approach demonstrated improved accuracy in detecting synthetic images across multiple domains.
2.4.3 Two-Stream Convolutional Network for Fake Content Detection by Yousaf: 2022